If n cells each of e.m.f ‘E’are connected in parallel,the e.m.f of the combination is
a)E/n  b)nE  c)n2/E  d)E
 

Grade:12

3 Answers

Siddharth Singh
9 years ago
Do you know the answer ? Cause I am not quite sure. 
S Ashish
9 years ago
In parallel combination of batteries, the resultant emf remains the same. Answer d)E
Siddharth Singh
9 years ago
Since
Eeq/Req=(E1/R1)+(E2/R2)......
So,
To find Eeq we first find req
i.e 
1/Req=1/R1+1/R2+....
common 1/R we have Req=R/n
Putting in formula. 
Eeq=(En/R)*R/n
We get E